OpenStudy (anonymous):

How do I find the midpoint? A(3, –5) and C(2, 9)

OpenStudy (anonymous):

Midpoint is average of coordinate: x = ( 3 +2 ) /2 = -- y = ( -5 + 9 ) /2 = ..